# GlobalGeoTree: A Multi-Granular Vision-Language Dataset for Global Tree Species Classification

### v1

@@ 4.1 GeoTreeCLIP model
To establish baseline performance on the GlobalGeoTree dataset, we developed GeoTreeCLIP, a vision-language model specifically designed for tree species classification. The core motivation for adopting the CLIP architecture lies in its ability to learn general and powerful representations through multimodal contrastive learning, which jointly trains on image-text pairs at scale. This approach not only aligns visual and textual modalities but also demonstrates strong transfer capabilities, particularly in zero-shot and few-shot scenarios [14]. Such capabilities are especially valuable given the continuously expanding species catalog and the practical limitations of obtaining exhaustive labeled data for all taxa. These qualities make the CLIP architecture particularly suitable for tree species classification, where leveraging rich textual descriptions (e.g., multi-level taxonomic labels) enhances image understanding and addresses challenges such as rare species representation and complex hierarchical structures.
@@ 4.3.1 Zero-shot evaluation
The experimental results reveal two critical patterns. First, the accuracy consistently declines as the taxonomic level becomes finer, reflecting the increasing complexity of distinguishing between closely related classes. This trend is evident across all models but is particularly severe for CLIP and RemoteCLIP, which exhibit minimal capability at the genus and species levels. In contrast, GeoTreeCLIP maintains stronger performance even at these fine-grained levels, likely due to its ability to learn and leverage the hierarchical relationships inherent in the taxonomic labels, as qualitatively supported by the feature embedding visualizations (see Appendix D and Figure 7). Second, the substantial performance gap between GeoTreeCLIP and the baseline models highlights the importance of domain-specific pretraining. GeoTreeCLIP’s ability to incorporate spatiotemporal and multispectral information enables it to address the unique challenges of tree species classification, which are beyond the scope of general-purpose models like CLIP and RemoteCLIP.
@@ Appendix C Model Performance on GlobalGeoTree-10kEval-300 and GlobalGeoTree-10kEval-900
Appendix D Qualitative Analysis: t-SNE Feature Embeddings Visualization
To qualitatively evaluate feature representations, we used t-SNE [49] to visualize image embeddings from samples in the GlobalGeoTree-10kEval-300 dataset, comparing GeoTreeCLIP with the original CLIP. A hierarchical analysis was conducted: first examining five randomly selected families, then five genera within Rosaceae, and finally five species within Prunus. Figure 7 shows GeoTreeCLIP (top row) consistently forming clearer, more distinct clusters compared to CLIP (bottom row). At the family level (left column), GeoTreeCLIP separates families like Hernandiaceae and Rosaceae better, while CLIP shows significant overlap. At the genus level (middle column), GeoTreeCLIP clusters genera such as Rhodotypos and Rubus distinctly, whereas CLIP merges them into a dense structure. At the species level (right column), GeoTreeCLIP demonstrates its ability to effectively distinguish between five Prunus species (P. avium, P. caroliniana, P. ilicifolia, P. laurocerasus, P. obtusata), whereas CLIP struggles to differentiate them. These results demonstrate that GeoTreeCLIP learns more semantically meaningful and discriminative representations across taxonomic levels, aligning with its superior quantitative performance.
Figure 7: t-SNE visualization of feature embeddings extracted by GeoTreeCLIP and Original CLIP across different taxonomic levels, using samples from GlobalGeoTree-10kEval-300.

### v2

@@ 4.1 GeoTreeCLIP model
To establish baseline performance on the GlobalGeoTree dataset, we developed GeoTreeCLIP, a vision-language model specifically designed for tree species classification. The core motivation for adopting the CLIP architecture lies in its ability to learn general and powerful representations through multimodal contrastive learning, which jointly trains on image-text pairs at scale. This approach not only aligns visual and textual modalities but also demonstrates strong transfer capabilities, particularly in zero-shot and few-shot scenarios [14]. Such capabilities are especially valuable given the continuously expanding species catalog and the practical limitations of obtaining exhaustive labeled data for all taxa. These qualities make the CLIP architecture particularly suitable for tree species classification, where leveraging rich textual descriptions (e.g., multi-level taxonomic labels) enhances image understanding and addresses challenges such as rare species representation and complex hierarchical structures. A detailed comparison illustrating the advantages of this contrastive learning approach over a traditional supervised learning paradigm is provided in Appendix D.
@@ 4.3.1 Zero-shot evaluation
The experimental results reveal two key patterns. First, accuracy consistently declines as the taxonomic level becomes finer, reflecting the growing challenge of distinguishing closely related classes. This trend is observed across all models but is particularly pronounced for CLIP and RemoteCLIP, which perform poorly at the genus and species levels. In contrast, GeoTreeCLIP demonstrates stronger performance at these fine-grained levels, likely due to its ability to learn and leverage the hierarchical relationships in taxonomic labels, as supported by feature embedding visualizations (see Appendix E and Figure 7). Second, the significant performance gap between GeoTreeCLIP and the baseline models underscores the importance of domain-specific pretraining. Moreover, its integration of spatiotemporal and multispectral information further enhances its ability as general-purpose models like CLIP and RemoteCLIP struggle to handle. Additional zero-shot benchmark results, including evaluations of SkyCLIP-50 [25] and CLIP-laion-RS, are provided in Appendix F.
@@ Appendix C Model Performance on GlobalGeoTree-10kEval-300 and GlobalGeoTree-10kEval-900
Appendix D Comparison with Supervised Learning Paradigm
To further contextualize the performance of our contrastive learning-based GeoTreeCLIP model, we conducted an additional experiment using a traditional supervised learning paradigm. This allows for a more direct comparison of learning objectives (contrastive vs. supervised) while keeping core architectural components and training settings as consistent as possible.
D.1 Supervised Model Architecture and Training
We designed a supervised model, termed SupervisedGeoTree, which retains the visual processing pathway of GeoTreeCLIP, including the VisualEncoder for Sentinel-2 time series and the AuxiliaryEncoder for environmental variables. The features from these two encoders are projected, normalized, and then fused via concatenation followed by a fusion layer, similar to the visual feature preparation in GeoTreeCLIP.
However, unlike GeoTreeCLIP, the SupervisedGeoTree model does not include a text encoder or employ a contrastive loss. Instead, the fused visual-auxiliary features are fed into four independent classification heads (fully connected layers), each dedicated to predicting labels for one of the hierarchical taxonomic levels: functional type (level0), family (level1_family), genus (level2_genus), and species (level3_species). The number of output neurons for each head corresponds to the number of unique classes at that respective taxonomic level in the GlobalGeoTree (4 for level0, 275 for family, 2,734 for genus, and 21,001 for species).
The model was trained on the GlobalGeoTree-6M dataset. The loss function employed was a sum of standard Cross-Entropy losses, calculated independently for each of the four taxonomic levels. The contributions of each level’s loss to the total loss were equally weighted. All other training hyperparameters, including the learning rate ($1\times 10^{-5}$), optimizer (AdamW), weight decay, number of epochs (25), warmup strategy (5 epochs), learning rate scheduler (Cosine Annealing with Warm Restarts), and batch size, were kept identical to those used for pretraining GeoTreeCLIP to ensure a fair comparison of the learning paradigms.
D.2 Zero-Shot Evaluation of the Supervised Model
After training on the GlobalGeoTree-6M dataset, the SupervisedGeoTree model was evaluated on the GlobalGeoTree-10kEval subset. Since this model is trained with fixed classification heads for the classes seen during training, its ability to perform "zero-shot" classification in the same sense as a CLIP-style model (i.e., classifying entirely new, unseen samples provided at test time) is inherently limited. However, for this comparison, we evaluate its performance on the classes within GlobalGeoTree-10kEval that were also part of the GlobalGeoTree-6M training vocabulary for each respective taxonomic head. If a class in GlobalGeoTree-10kEval was not in the training vocabulary for a specific head, it cannot be correctly predicted by that head.
Notably, the GlobalGeoTree-6M dataset was designed to retain nearly all tree species categories, as our goal was to pretrain a model capable of classifying the full spectrum of tree species. Consequently, the zero-shot evaluation here can be seen as measuring the model’s “zero-shot” transfer capability on unseen datasets, aligning with the concept of "in-domain" zero-shot classification defined in [25].
Table 10: Zero-shot evaluation on GlobalGeoTree-10kEval. SupervisedGeoTree is evaluated on classes within its training vocabulary. Results are mean accuracy (%) $\pm$ standard deviation (%) over 5 runs.
Taxon. CLIP RemoteCLIP SupervisedGeoTree GeoTreeCLIP
Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
Family 10.80 $\pm$ 0.03 25.32 $\pm$ 0.05 1.11 $\pm$ 0.01 10.55 $\pm$ 0.04 9.55 $\pm$ 0.10 27.41 $\pm$ 0.07 20.99 $\pm$ 0.28 56.88 $\pm$ 0.42
Genus 1.09 $\pm$ 0.01 9.34 $\pm$ 0.01 1.11 $\pm$ 0.01 6.25 $\pm$ 0.02 1.19 $\pm$ 0.09 8.18 $\pm$ 0.14 18.39 $\pm$ 0.26 50.98 $\pm$ 0.41
Species 1.09 $\pm$ 0.01 7.02 $\pm$ 0.02 1.11 $\pm$ 0.01 6.25 $\pm$ 0.02 0.00 $\pm$ 0.00 0.28 $\pm$ 0.02 16.71 $\pm$ 0.25 47.52 $\pm$ 0.37
The results in Table 10 indicate that a consistent trend across all models is the decline in classification accuracy as the taxonomic level becomes finer (from family to species). This reflects the inherent increase in difficulty when distinguishing between more closely related taxa. However, the extent of this performance degradation varies significantly between models.
The SupervisedGeoTree model, which employs traditional supervised classification heads for each taxonomic level, achieves reasonable accuracy at the family level (9.55% top-1). However, its performance drops sharply for genus (1.19% top-1) and becomes negligible at the species level (0.00% top-1). This drastic decline underscores the challenge of fine-grained classification when relying solely on visual and auxiliary features without leveraging the semantic relationships embedded in textual taxonomic labels, and the inherent limitation of generalizing to a large number of specific classes in a purely supervised manner.
In stark contrast, our proposed GeoTreeCLIP model demonstrates substantial improvements over all other baselines across every taxonomic level. At the family level, GeoTreeCLIP achieves a top-1 accuracy of 20.99%, more than doubling the performance of CLIP and significantly outperforming SupervisedGeoTree. This advantage is even more pronounced at the finer-grained levels: GeoTreeCLIP obtains 18.39% top-1 accuracy for genus and 16.71% for species identification. These results strongly indicate the power of contrastive vision-language learning to align visual features with rich, hierarchical taxonomic text labels. This capability allows GeoTreeCLIP to learn more nuanced and generalizable representations, leading to its superior performance in this challenging zero-shot evaluation. The significant gap, especially compared to SupervisedGeoTree at the species level, highlights the efficacy of the contrastive learning paradigm for handling large, structured label spaces.
Appendix E Qualitative Analysis: t-SNE Feature Embeddings Visualization
To qualitatively assess and compare the learned feature representations from different modeling paradigms, we performed t-SNE visualizations [49]. We used zero-shot image embeddings (or the final fused visual-auxiliary features for SupervisedGeoTree before the classification heads) extracted from a subset of the GlobalGeoTree-10kEval-300 dataset. This analysis includes our proposed GeoTreeCLIP, the original CLIP pretrained by OpenAI, and the SupervisedGeoTree model. Given the large number of classes and the hierarchical nature of the labels, we adopted a selective visualization strategy: first examining embeddings at the family level for five randomly selected families (Anacardiaceae, Berberidaceae, Cactaceae, Hernandiaceae, and Rosaceae); then focusing on five genera within the Rosaceae family (Prunus, Pyrus, Rhodotypos, Rubus, and Spiraea); and finally, visualizing five species within the Prunus genus (Prunus avium, Prunus caroliniana, Prunus ilicifolia, Prunus laurocerasus, and Prunus obtusata).
The comparative t-SNE visualizations are presented in Figure 7. Across all three taxonomic levels (Family, Genus, and Species, shown as columns), GeoTreeCLIP (top row) consistently demonstrates the most effective separation and formation of distinct clusters. At the family level (left column), GeoTreeCLIP clearly distinguishes between the selected families. The original CLIP (middle row) exhibits considerable overlap, particularly for the diverse Rosaceae family. SupervisedGeoTree (bottom row) shows some separation but less defined clusters compared to GeoTreeCLIP, with Rosaceae still forming a very broad distribution.
Figure 7: t-SNE visualization of feature embeddings from GlobalGeoTree-10kEval-300 at different taxonomic levels, comparing GeoTreeCLIP (top row), original CLIP (middle row), and SupervisedGeoTree (bottom row). Columns from left to right represent visualizations at the Family level (selected: Anacardiaceae, Berberidaceae, Cactaceae, Hernandiaceae, Rosaceae), Genus level (selected within Rosaceae: Prunus, Pyrus, Rhodotypos, Rubus, Spiraea), and Species level (selected within Prunus: Prunus avium, Prunus caroliniana, Prunus ilicifolia, Prunus laurocerasus, Prunus obtusata).
This pattern of superior clustering by GeoTreeCLIP continues at the genus level within Rosaceae (middle column). GeoTreeCLIP forms relatively distinct groups for genera like Prunus, Pyrus, and Rhodotypos. Both CLIP and SupervisedGeoTree struggle more, with CLIP showing a highly condensed and overlapping structure, while SupervisedGeoTree offers some separation but with less clarity than GeoTreeCLIP. The most striking difference is observed at the species level within the Prunus genus (right column). GeoTreeCLIP achieves remarkable separation, forming visually distinct clusters for each of the five Prunus species. In contrast, both original CLIP and SupervisedGeoTree largely fail to differentiate these closely related species, with their embeddings heavily intermingled.
These visualizations qualitatively affirm that GeoTreeCLIP, through its contrastive vision-language learning approach tailored with hierarchical taxonomic information and domain-specific data, learns more semantically meaningful and discriminative representations across all taxonomic ranks. This aligns with its superior quantitative performance in classification tasks compared to both general-domain VLMs and a traditional supervised approach.
Appendix F Additional Baselines Evaluated on GlobalGeoTree-10kEval
To further contextualize GeoTreeCLIP’s performance, we extended our zero-shot evaluation on the GlobalGeoTree-10kEval subset to include two additional publicly available vision-language models: SkyCLIP-50 [25] and CLIP-laion-RS, a CLIP model pretrained on the remote sensing subset of LAION-2B [50]. These models were evaluated on the GlobalGeoTree-10kEval subset under the same zero-shot protocol used for original CLIP and RemoteCLIP (features extracted from individual monthly images, probabilities averaged). For ease of comparison, their performance alongside our GeoTreeCLIP is presented in Table 11.
Table 11: Zero-shot evaluation of SkyCLIP-50, CLIP-laion-RS, and GeoTreeCLIP on GlobalGeoTree-10kEval. Results are presented as mean accuracy (%) $\pm$ standard deviation (%) over 5 runs.
Taxon. SkyCLIP-50 CLIP-laion-RS GeoTreeCLIP
Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
Family 2.33 $\pm$ 0.01 18.40 $\pm$ 0.08 1.15 $\pm$ 0.01 17.83 $\pm$ 0.16 20.99 $\pm$ 0.28 56.88 $\pm$ 0.42
Genus 1.10 $\pm$ 0.03 6.48 $\pm$ 0.04 1.12 $\pm$ 0.01 7.33 $\pm$ 0.03 18.39 $\pm$ 0.26 50.98 $\pm$ 0.41
Species 1.10 $\pm$ 0.03 6.36 $\pm$ 0.04 1.12 $\pm$ 0.01 7.27 $\pm$ 0.02 16.71 $\pm$ 0.25 47.52 $\pm$ 0.37
The results in Table 11 show that both SkyCLIP-50 and CLIP-laion-RS, despite their pretraining on remote sensing imagery, achieve zero-shot accuracies on GlobalGeoTree-10kEval that are substantially lower than our GeoTreeCLIP. For instance, at the species level, SkyCLIP-50 obtains a top-1 accuracy of 1.10% and CLIP-LAION-RS achieves 1.12%, in contrast to GeoTreeCLIP’s 16.71%. As indicated in the main text for similar baseline models (original CLIP, RemoteCLIP), such performance can partly be attributed to their design, which is often optimized for RGB data and lacks effective handling of time-series information or small-patch classification crucial for tree species identification. These limitations further emphasize the value of the GlobalGeoTree benchmark and the effectiveness of our tailored GeoTreeCLIP approach in advancing global tree species classification research.
The comparison highlights that general remote sensing pretraining alone is insufficient for the nuanced task of global fine-grained tree species identification. The domain-specific dataset characteristics, multimodal input integration (including time-series and auxiliary data), and the tailored contrastive learning approach of GeoTreeCLIP appear critical for achieving strong performance on this challenging benchmark.